Every student at this University is classified as:

- in-state or out-of-state

- male or female

- living on-campus or living off-campus

Of the students 3,000 are in-state, 4,600 are male, and 7,000 live on-campus. Also, 1,320 are in-state males and 3,010 are males that live on campus and 1,400 are in-state and live on campus. Finally, 600 are in-state males that live on-campus.

If currently there are 10,000 students at this University, calculate the number of female students that are in-state and live off-campus.
Answer #1

Given

S=3000

S M C =600

S M = 1320

S M C' =S M - S M C = 1320 -600 = 720

S C = 1400

S F C = S C - S M C = 1400-600 = 800

S F C' = S- S M C - S M C' -S F C = 3000- 600- 720- 800 = 880

Thus number of female students that are in state and live off campus = 880